OpenStudy (anonymous):

If the equation y=2^x is graphed, which of the following values of x would produce a point closest to the x-axis? A) 1/4 B) 3/4 C) 5/3 D)8/3

OpenStudy (anonymous):

do you know how the graph of y=2^x looks like?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about answering this... if a point was very close to the x-axis, what would you be able to tell me about the y-coordinate of that point?

OpenStudy (anonymous):

y=0 ?

OpenStudy (anonymous):

well, yes... the point's y-coordinate would be very close to 0. so which of those choices will give you the smallest y-value?

OpenStudy (anonymous):

1/4?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

in other words, which is the smallest value? \(\large 2^{\frac{1}{4}} \) OR \(\large 2^{\frac{3}{4}} \) OR \(\large 2^{\frac{5}{3}} \) OR \(\large 2^{\frac{8}{3}} \) ????

OpenStudy (anonymous):

2 and 1/4

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@dpaInc ..?

OpenStudy (anonymous):

is it to the power of? y=2^x so would it be 2^1/4

OpenStudy (anonymous):

those aren't fractions... \(\large 2^{\frac{1}{4}} \)= "two raised to the 1/4 power"

OpenStudy (anonymous):

oh i see... so

OpenStudy (anonymous):

so is it A because it is the smallest value, getting the closest to the x-axis?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

??

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

yes... :)

OpenStudy (anonymous):

a is correct?